# The Nuclear Star Cluster of M 74: a fossil record of the very early stages of a star-forming galaxy

### v1

@@
(December 3, 2025)
Nuclear star clusters (NSC) are dense and compact stellar systems, of sizes of few parsecs, located at galactic centers. Their properties and formation mechanisms seem to be tightly linked to the evolution of the host galaxy, with potentially different formation channels for late- and early-type galaxies (respectively, LTGs and ETGs). While most observations target ETGs, here we focus on the NSC in M 74 (NGC 628), a relatively massive, gas-rich and star-forming spiral galaxy, part of the PHANGS survey. We analyzed the central arcmin of the PHANGS-MUSE mosaic, in which the NSC is not spatially resolved. We performed a two-dimensional spectro-photometric decomposition of the MUSE cube, employing a modified version of the C2D code, to disentangle the NSC from the host galaxy. Here we used three components: a bulge, a disk and a NSC approximated to the point spread function (PSF), obtaining three data cubes, one for each component. This allowed us to extract separately the age, metallicity and [Mg/Fe] abundance for the NSC and the host galaxy. Our results show a very old and metal-poor NSC, in contrast to the surrounding regions. While similar properties were found in NSCs hosted by galaxies of different masses and/or morphological types from M 74, they are somewhat unexpected for a relatively massive star-forming spiral galaxy. The spatially resolved stellar populations of the host galaxy display much younger (light-weighted) ages and higher metallicities, especially in the central region (${\sim}500$ pc) surrounding the NSC. This suggests that this NSC formed a long time ago, and evolved passively until today, without any further growth. Our results show that the NSC was not involved in the active recent star-formation history of its host galaxy.
@@ 1 Introduction
In the central regions of galaxies, massive black holes often coexist with dense stellar systems known as nuclear star clusters (NSCs). NSCs have sizes of a few parsecs and masses between $10^{4}$ and a few times $10^{8}$ M(e.g., Cote2006; Georgiev2016; SanchezJanssen2019; Hoyer2023a). Their properties trace the dynamical state and evolution of the central regions of galaxies and are key to revealing not only their formation and growth but also the mechanisms funneling stars and gas to the center, leading to the seeding and growth of central black holes and to the appearance of active galactic nuclei (AGNs, see Neumayer2020 for a review). NSCs can form via two main channels: star-cluster inspiral to the center due to dynamical friction (e.g., Tremaine1975; Capuzzo-Dolcetta1993; Oh2000; Lotz2001; Arcasedda2014) and in-situ star formation after gas inflow (e.g., Silk1987; Mihos1994; Milosavljevic2004; Bekki2006; Bekki2007). These two formation channels are not mutually exclusive, as migrating star clusters can be gas-rich (Guillard2016), and combinations of in-situ formation and cluster merging have often been necessary to explain the observed properties (Fahrion2019; Fahrion2021; Fahrion2022a; Fahrion2022b; Fahrion2024).
@@
More than 60% of galaxies with stellar masses between $10^{8}$ and $10^{10}$ M host an NSC, with a peak nucleation fraction (${\sim}90$%) near $10^{9}$ M (SanchezJanssen2019; Hoyer2021). This fraction drops drastically for lower masses. For higher masses, a similar drop is observed for early-type galaxies (ETGs) hosting NSCs, while massive late-type galaxies (LTGs) seem to maintain a similar high fraction (Neumayer2020; Ashok2023). This trend, together with NSC properties such as metallicity, suggests a transition in the dominant NSC formation channel at about $10^{9}$ M(e.g. Lyu2025): many low-mass galaxies host metal-poor NSCs, which have similar properties to globular clusters and are traditionally associated with cluster inspiral (e.g., Fahrion2020). NSCs are systematically more metal rich at higher galaxy masses, showing significant contributions from younger stellar populations, often ongoing or recent star formation (see, e.g., Fig. 9 in the review from Neumayer2020, and Fahrion2021; Fahrion2022a; Fahrion2022b). However, metal-rich NSCs may also form from the inspiral of young massive star clusters, complicating more the interpretation of observations (Paudel2020; Fahrion2024).
@@
Most evidence of the trends described above comes from NSCs hosted by ETGs, while it needs to be further investigated in LTGs. Several kinematic and stellar-population studies revealed morphological dependencies. Pinna2021 analyzed the spatially resolved kinematics of eleven NSCs. While rotation is ubiquitous in the full sample, NSCs in LTGs are typically rotation-dominated and associated with central gas inflow, whereas those in ETGs show slower rotation, more complexity and signatures of past mergers (especially at lower NSC masses). NSCs in very LTGs (Scd or later types) are on average young, but made up of stellar populations of different ages, with prolonged and recurrent star formation until the present day (Walcher2005; Walcher2006; Rossa2006; Kacharov2018). NSCs in LTGs also showed a diversity in their ages and metallicities (Kacharov2018). NSCs are on average older in earlier types spirals (${\sim}1-2$ Gyr) and younger, less massive, and more metal poor in the latest types (about a couple of hundred Myr, Rossa2006; Walcher2006). Combined trends with galaxy mass and morphological types support a link between NSC formation and the host-galaxy evolution.
@@
While simulations have reproduced aspects of both formation channels, systematic theoretical studies across galaxy types are still lacking. On the one hand, individual parsec-scale resolution $N$-body simulations were essential to show that star cluster inspiral can easily lead to the observed properties of NSCs (e.g., Mastrobuono2014; Arcasedda2014; ArcaSedda2020; Mastrobuono2023; Mastrobuono2025). On the other hand, other simulations have shown that in-situ formation is needed to explain the complexity in the kinematics and stellar population properties of stars in the center, e.g., of the Milky Way (e.g., Mapelli2012a; Mastrobuono2019). However, in numerical simulations, larger galaxy samples and a cosmological context are needed to assess the impact of internal and external galaxy-evolution processes in the formation of nuclear structures, while the study of NSCs requires very high spatial and mass resolution, posing a serious computational challenge. Cosmological simulations of galaxies at redshift $z{\sim}1.5$ by Brown2018 showed the presence of central clusters (made up of only few stellar particles), with extended star formation histories (SFHs) leading to large spreads in ages and metallicities, and global higher metallicities than their host galaxy. A recent study using the Engineering Dwarfs at Galaxy Formation’s Edge (EDGE) simulations (Agertz2020) traced back the emergence of an NSC in four models of low-mass dwarf galaxies (with stellar masses between $10^{6}$ and $10^{7}$ M, Gray2025). The NSC forms at redshift $z{\sim}2$ during a starburst triggered by a major merger, but also contains a previously formed population of stars.
@@
Despite this progress, NSC formation remains unclear, and the relative importance of different growth mechanisms is still debated. Spatially resolving NSCs with spectroscopic observations is currently feasible only in nearby bright galaxies. Although integral-field spectroscopy (IFS) enables comparisons between the NSC and its host galaxy, in most cases, NSCs are still unresolved, and we are limited to their point-spread-function (PSF) integrated spectra (e.g., Fahrion2021; Fahrion2022b). Yet, due to the small size of these objects, these studies are challenging and IFS samples are mostly restricted to ETGs (see above), mainly because lower surface brightness in LTGs requires long integration times. While NSCs in some dwarf LTGs were recently covered (Fahrion2022b), additional work is needed on massive spirals.
@@
Accurate measurements further require disentangling NSC light from that of the host galaxy, to limit the contamination from other internal structures overlapped in the line of sight (LOS). This problem has been approached in different ways in the past. Johnston2020 used BUDDI (Bulge–Disc Decomposition of IFU data, Johnston2017) to model the light of the NSC and the host galaxy in two dimensions, for each wavelength, and separate one spectrum for each component. They modeled the NSC as a PSF profile and the host galaxy as a combination of one to four Sérsic profiles. This ensured a minimal contamination of the NSC spectrum from the host galaxy. Similarly, Fahrion2021 simply obtained the NSC spectrum by subtracting a representative spectrum of the host galaxy from a central PSF-weighted spectrum. The host-galaxy spectrum was obtained by multiplying a spectrum extracted from an elliptical annulus around the central region (at 160200160-200pc from the center) by a scaling factor, calculated by modeling images for the NSC (as a PSF) and the host galaxy (the latter, modeled as a combination of a bulge and a disk). In this way, they preserved the spatial gradients.
@@
In this work, we present a new spectro-photometric decomposition method that separates an IFS data cube into multiple morphological components, enabling the extraction of NSC properties while mapping its surroundings. We applied this method to the central region of the massive spiral galaxy M 74 (NGC 628) and analyzed its stellar populations. The paper is structured as follows. In Sect. 2, we describe the data set. Section 3 summarizes the known properties of the galaxy M 74 and its NSC. In Sect. 4, we describe the methods used for the spectro-photometric decomposition of the NSC from the host galaxy, and for the stellar-population analysis. The analysis of the original IFS data cube is presented in Appendix A. Section 5 describes our results, which are discussed in Sect. 6. Our conclusions are outlined in Sect. 7.
@@ 2 Observations
M 74 was observed with 12 pointings covering the full disk (Kreckel2016; Kreckel2017; Kreckel2018; Program IDs: 094.C-0623, 095.C-0473, 098.C-0484; P.I.s: K. Kreckel and G. Blanc). An image reconstructed from the full PHANGS-MUSE mosaic is shown in Fig. 9 of Emsellem2022 (Emsellem2022, see also Fig. 1 in Kreckel2018). Each pointing was observed with three exposures of 845 or 990 seconds (depending on the pointing), rotated 90 with respect to each other. Sky exposures were taken in between the object exposures. The data reduction was carried out using the dedicated implementation of the MUSE data reduction recipes (Weilbacher2020) into the wrapper pymusepipe [^2] https://pypi.org/project/pymusepipe/. The data reduction process included corrections from bias, flat field, wavelength calibration, line-spread-function derivation, geometry and astrometry corrections, illumination correction through twilight sky flats, sky subtraction, alignment and mosaicking. The full data reduction process was detailed by Emsellem2022. The reduced mosaic is publicly available in the ESO Archive [^3] https://archive.eso.org/scienceportal/home?data_collection=PHANGS.
Because individual pointings were observed at different times and thus with different atmospheric conditions, the PSF varied between each other and was modeled for different pointings using a circular Moffat function, which described it well (Fusco2020). The Moffat FWHM and power index were measured for each pointing using a cross-convolution method (by comparing with a reference image) and were assumed constant within each MUSE FoV (Emsellem2022). In this work, we adopted a PSF FWHM of 3 MUSE spatial pixels (3 spaxels, 0.6 arcsec), as this is the approximate FWHM measured in the pointings contributing to the cropped central data cube. The power index estimate was 2.8 for all pointings. The MUSE PSF also varies with wavelength, and it was measured at a reference wavelength of 6483.58 Å. We tested that a variation of up to $\pm 0.2$ arcsec (one pixel) in the FWHM did not affect the results (Sect. 4.1).
@@ 3 M 74 (NGC 628)
M 74 (NGC 628, or the ”Phantom Galaxy”) is a grand-design non-barred spiral galaxy classified as an SA(s)c (Buta2015). It lies at a distance of 9.84 Mpc and has a stellar mass of $2.2\times 10^{10}$ M(Leroy2021). Its nearly face-on inclination (${\sim}8.9\mbox{${}^{\circ}$}$, Lang2020) is optimal to disentangle properties of the stellar nucleus from the host galaxy. Morphologically, M 74 consists of a faint bulge and a dominating disk, with no obvious bar (Salo2015; Querejeta2021; Stuber2023)), though remnants of a past bar and a possible nuclear ring have been debated (e.g., SanchezBlazquez2014b).
@@
M 74 is a star-forming galaxy, with a star-formation rate of $\log_{10}\,(\textrm{SFR}/[\textrm{M}_{\odot}/\textrm{yr}])=0.24$, which places it above the star-formation main sequence of local galaxies (Leroy2021; Kim2023). Its star formation and interstellar medium have been extensively studied at high spatial resolution by the PHANGS team using a combination of MUSE, ALMA, HST, and JWST. Massive star-forming regions are observed both in the spiral arm and interarm regions (Kreckel2016; Williams2022). On large scales, star formation mostly follows the molecular-gas structure, while offsets at the scale of molecular clouds imply longer gas depletion times than in the Milky Way (Kreckel2018). HST and JWST images revealed with unprecedented detail a complex web of dust and gas filaments, ranging from compact molecular-cloud-sized structures to faint arm-tracing features (Thilker2023), along with stellar-feedback-driven bubbles and voids (Barnes2023; Watkins2023). A remarkable cavity of a size of about $200\times 400$ pc2, with no ionized or molecular gas, and no dust, was found in the center of the galaxy (Kreckel2018; Herrera2020; Leroy2021; Stuber2023; Hoyer2023b).
@@
M 74’s spatially resolved stellar populations show radial trends. At large scale, SanchezBlazquez2014b found predominantly old ages (${\sim}10$ Gyr) beyond ${\sim}6$ kpc, with a negative gradient consistent with an inside-out disk formation. The SFH reveals a dominant ${\sim}13$ Gyr component, plus a minor recent ${\sim}1$ Gyr component. The metallicities are slightly subsolar with a relatively flat gradient within ${\sim}3$ kpc, which becomes negative outwards. A more recent PHANGS-MUSE analysis (Pessa2023), confirmed these trends at ${\sim}100$ pc resolution, although with average light-weighted and mass-weighted ages in the central kpc, respectively, of about 2 and 9 Gyr, and the youngest stellar populations along the spiral arms. Metallicity gradients are negative in the central kpc and flatter (slightly positive) in the outskirts, again supporting an inside-out assembly scenario. JWST resolved-star analyses further confirmed the negative metallicity gradient (Ck2025). Finally, the star-cluster analysis from HST data has shown a distribution of ages ranging from a few million years, for cluster masses between $10^{3}$ and $10^{5}$ M, to more than 10 billion years above $10^{5}$ M(up to $10^{7}$ M, Whitmore2023).
@@
M 74 is the brightest member of the M 74 Group, in which dwarf satellite galaxies (with stellar masses $\rm 10^{4}\,M_{\sun}<M_{*}<10^{6}$ M) identified by Davis2021 show signs of recent quenching. M 74’s HI disk shows a complex structure with a warp, high velocity complexes, and an asymmetric outer tail (Kamphuis1992). These features might be the result of one or more past accretion events.

### v2

@@
Nuclear star clusters (NSC) are dense and compact stellar systems, of sizes of few parsecs, located at galactic centers. Their properties and formation mechanisms seem to be tightly linked to the evolution of the host galaxy, with potentially different formation channels for late- and early-type galaxies (respectively, LTGs and ETGs). While most observations target ETGs, here we focus on the NSC in M 74 (NGC 628), a relatively massive, gas-rich and star-forming spiral galaxy, part of the PHANGS survey. We analyzed the central arcmin of the PHANGS-MUSE mosaic, in which the NSC is not spatially resolved. We analyzed the NSC stellar populations in a point spread function (PSF) aperture, and compared it to the host galaxy. Within the PSF size, the NSC is contaminated by the host-galaxy light. We performed a two-dimensional spectro-photometric decomposition of the MUSE cube, employing a modified version of the C2D code, to disentangle the NSC from its host. This method provided different data cubes for the NSC and the host galaxy, allowing for both their comparison in a PSF aperture, and the spatially resolved analysis of the host. Our results show a very old and metal-poor NSC, in contrast to the surrounding regions. While similar properties were found in NSCs hosted by galaxies of different masses and/or morphological types from M 74, they are somewhat unexpected for a relatively massive star-forming spiral galaxy. The spatially resolved stellar populations of the host galaxy display much younger (light-weighted) ages and higher metallicities, especially in the central region (${\sim}500$ pc) surrounding the NSC. This suggests that this NSC formed a long time ago, and evolved passively until today, without any further growth. Our results show that the NSC was not involved in the active recent star-formation history of its host galaxy.
@@ 1 Introduction
In the central regions of galaxies, massive black holes often coexist with dense stellar systems known as nuclear star clusters (NSCs). NSCs have sizes of a few parsecs and masses between $10^{4}$ and a few times $10^{8}$ M(e.g., Côté et al. 2006; Georgiev et al. 2016; Sánchez-Janssen et al. 2019; Hoyer et al. 2023a). Their properties trace the dynamical state and evolution of the central regions of galaxies and are key to revealing not only their formation and growth but also the mechanisms funneling stars and gas to the center, leading to the seeding and growth of central black holes (see Neumayer et al. 2020 for a review). NSCs can form via two main channels: star-cluster inspiral to the center due to dynamical friction (e.g., Tremaine et al. 1975; Capuzzo-Dolcetta 1993; Oh and Lin 2000; Lotz et al. 2001; Arca-Sedda and Capuzzo-Dolcetta 2014) and in-situ star formation after gas inflow (e.g., Silk et al. 1987; Mihos and Hernquist 1994; Milosavljević 2004; Bekki et al. 2006; Bekki 2007). These two formation channels are not mutually exclusive, as migrating star clusters can be gas-rich (Guillard et al. 2016), and combinations of in-situ formation and cluster merging have often been necessary to explain the observed properties (Fahrion et al. 2019, 2021, 2022b, 2022a, 2024).
@@
More than 60% of galaxies with stellar masses between $10^{8}$ and $10^{10}$ M host an NSC, with a peak nucleation fraction (${\sim}90$%) near $10^{9}$ M (Sánchez-Janssen et al. 2019; Hoyer et al. 2021). This fraction drops drastically for lower masses. For higher masses, a similar drop is observed for early-type galaxies (ETGs) hosting NSCs, while massive late-type galaxies (LTGs) seem to maintain a similar high fraction (Neumayer et al. 2020; Ashok et al. 2023). This trend, together with NSC properties such as metallicity, suggests a transition in the dominant NSC formation channel at about $10^{9}$ M(e.g. Lyu et al. 2025): many low-mass galaxies host metal-poor NSCs, which have similar properties to globular clusters and are traditionally associated with cluster inspiral (e.g., Fahrion et al. 2020). NSCs are systematically more metal rich at higher galaxy masses, showing significant contributions from younger stellar populations, often ongoing or recent star formation (see, e.g., Fig. 9 in the review from Neumayer et al. 2020, and Fahrion et al. 2021, 2022b, 2022a). However, metal-rich NSCs may also form from the inspiral of young massive star clusters, complicating more the interpretation of observations (Paudel and Yoon 2020; Fahrion et al. 2024).
@@
Most evidence of the trends described above comes from NSCs hosted by ETGs, while it needs to be further investigated in LTGs. Several kinematic and stellar-population studies revealed morphological dependencies. Pinna et al. (2021) analyzed the spatially resolved kinematics of eleven NSCs. While rotation is ubiquitous in the full sample, NSCs in LTGs are typically rotation-dominated and associated with central gas inflow, whereas those in ETGs show slower rotation, more complexity and signatures of past mergers (especially at lower NSC masses). NSCs in very LTGs (Scd or later types) are on average young, but made up of stellar populations of different ages, with prolonged and recurrent star formation until the present day (Walcher et al. 2005, 2006; Rossa et al. 2006; Kacharov et al. 2018). NSCs in LTGs also showed a diversity in their ages and metallicities (Kacharov et al. 2018). Despite the scatter, NSCs follow an average trend being in general older in earlier types spirals (${\sim}1-2$ Gyr) and younger, less massive, and more metal poor in the latest types (about a couple of hundred Myr, Rossa et al. 2006; Walcher et al. 2006). Combined trends with galaxy mass and morphological types support a link between NSC formation and the host-galaxy evolution.
@@
While simulations have reproduced aspects of both formation channels, systematic theoretical studies across galaxy types are still lacking. On the one hand, individual parsec-scale resolution $N$-body simulations were essential to show that star cluster inspiral can easily lead to the observed properties of NSCs (e.g., Mastrobuono-Battisti et al. 2014; Arca-Sedda and Capuzzo-Dolcetta 2014; Arca Sedda et al. 2020; Mastrobuono-Battisti et al. 2023, 2025). On the other hand, other simulations have shown that in-situ formation is needed to explain the complexity in the kinematics and stellar population properties of stars in the center, e.g., of the Milky Way (e.g., Mapelli et al. 2012; Mastrobuono-Battisti et al. 2019). However, in numerical simulations, larger galaxy samples and a cosmological context are needed to assess the impact of internal and external galaxy-evolution processes in the formation of nuclear structures, while the study of NSCs requires very high spatial and mass resolution, posing a serious computational challenge. Cosmological simulations of galaxies at redshift $z{\sim}1.5$ by Brown et al. (2018) showed the presence of central clusters (made up of only few stellar particles), with extended star formation histories (SFHs) leading to large spreads in ages and metallicities, and global higher metallicities than their host galaxy. A recent study using the Engineering Dwarfs at Galaxy Formation’s Edge (EDGE) simulations (Agertz et al. 2020) traced back the emergence of an NSC in four models of low-mass dwarf galaxies (with stellar masses between $10^{6}$ and $10^{7}$ M, Gray et al. 2025). The NSC forms at redshift $z{\sim}2$ during a starburst triggered by a major merger, but also contains a previously formed population of stars.
@@
Despite this progress, NSC formation remains unclear, and the relative importance of different growth mechanisms is still debated. Spatially resolving NSCs with spectroscopic observations is currently feasible only in nearby bright galaxies. Although integral-field spectroscopy (IFS) enables comparisons between the NSC and its host galaxy, in most cases, NSCs are still unresolved, and we are limited to their point-spread-function (PSF) integrated spectra (e.g., Fahrion et al. 2021, 2022a). Yet, due to the small size of these objects, these studies are challenging and IFS samples are mostly restricted to ETGs (see above), mainly because lower surface brightness in LTGs requires long integration times. While NSCs in some dwarf LTGs were recently covered (Fahrion et al. 2022a), additional work is needed on massive spirals.
@@
Accurate measurements further require disentangling NSC light from that of the host galaxy, to limit the contamination from other internal structures overlapped in the line of sight (LOS). This problem has been approached in different ways in the past. Johnston et al. (2020) used BUDDI (Bulge–Disc Decomposition of IFU data, Johnston et al. 2017) to model the light of the NSC and the host galaxy in two dimensions, for each wavelength, and separate one spectrum for each component. They modeled the NSC as a PSF profile and the host galaxy as a combination of one to four Sérsic profiles. This ensured a minimal contamination of the NSC spectrum from the host galaxy. Fahrion et al. (2021) simply obtained the NSC spectrum by subtracting a representative spectrum of the host galaxy from a central PSF-weighted spectrum. The host-galaxy spectrum was obtained by multiplying a spectrum extracted from an elliptical annulus around the central region (at 160200160-200pc from the center) by a scaling factor, calculated by modeling images for the NSC (as a PSF) and the host galaxy (the latter, modeled as a combination of a bulge and a disk).
@@
In this work, we present a new version of the spectro-photometric decomposition method C2D that enables the extraction of NSC properties while mapping its surroundings. C2D (Méndez-Abreu et al. 2019) separates an IFS data cube into multiple morphological components, delivering a data cube for each one of them. It uses a very similar approach to BUDDI (see above), with the difference that the latter delivers one spectrum for each morphological component, instead of a full separated data cube as C2D does. This allows us to spatially resolve stellar populations in the NSC surroundings, helping understand the NSC evolution in the context of its host galaxy. We applied this method to the central region of the massive spiral galaxy M 74 (NGC 628) and analyzed its stellar populations. The paper is structured as follows. In Sect. 2, we describe the data set. Section 3 summarizes the known properties of the galaxy M 74 and its NSC. In Sect. 4, we describe the methods used for the spectro-photometric decomposition of the NSC from the host galaxy, and for the stellar-population analysis. The analysis of the non-decomposed IFS data cube is presented in Appendix A. Section 5 describes our results, which are discussed in Sect. 6. Our conclusions are outlined in Sect. 7.
@@ 2 Observations
M 74 was observed with 12 pointings covering the full disk (Kreckel et al. 2016, 2017, 2018; Program IDs: 094.C-0623, 095.C-0473, 098.C-0484; P.I.s: K. Kreckel and G. Blanc). An image reconstructed from the full PHANGS-MUSE mosaic is shown in Fig. 9 of Emsellem et al. (2022, see also Fig. 1 in Kreckel et al. 2018). Each pointing was observed with three exposures of 845 or 990 seconds (depending on the pointing), rotated 90 with respect to each other. Sky exposures were taken in between the object exposures. The data reduction was carried out using the dedicated implementation of the MUSE data reduction recipes (Weilbacher et al. 2020) into the wrapper pymusepipe [^2] https://pypi.org/project/pymusepipe/. The data reduction process included corrections from bias, flat field, wavelength calibration, line-spread-function derivation, geometry and astrometry corrections, illumination correction through twilight sky flats, sky subtraction, alignment and mosaicking. The full data reduction process was detailed by Emsellem et al. (2022). The reduced mosaic is publicly available in the ESO Archive [^3] https://archive.eso.org/scienceportal/home?data_collection=PHANGS.
Because individual pointings were observed at different times and thus with different atmospheric conditions, the PSF varied between each other and was modeled for the PHANGS-MUSE survey for different pointings (Emsellem et al. 2022), using a circular Moffat function, which described it well (Fusco et al. 2020). The Moffat FWHM and power index were measured for each pointing using a cross-convolution method (by comparing with a reference image) and were assumed constant within each MUSE FoV (Emsellem et al. 2022). In this work, we used a PSF FWHM of 3 MUSE spatial pixels (3 spaxels, 0.6 arcsec) for the NSC decomposition, which approximates the average FWHM measured in the four pointings contributing to the central arcmin. The power index estimate was 2.8 for all pointings. The MUSE PSF also varies with wavelength, and it was measured by Emsellem et al. (2022) at a reference wavelength of 6483.58 Å. We tested that a variation of up to $\pm 0.2$ arcsec (one pixel) in the FWHM did not affect the results (Sect. 4.1).
@@ 3 M 74 (NGC 628)
M 74 (NGC 628, or the ”Phantom Galaxy”) is a grand-design non-barred spiral galaxy classified as an SA(s)c (Buta et al. 2015). It lies at a distance of 9.84 Mpc and has a stellar mass of $2.2\times 10^{10}$ M(Leroy et al. 2021). Its nearly face-on inclination (${\sim}8.9\mbox{${}^{\circ}$}$, Lang et al. 2020) is optimal to disentangle properties of the stellar nucleus from the host galaxy. Morphologically, M 74 consists of a faint bulge and a dominating disk, with no obvious bar (Salo et al. 2015; Querejeta et al. 2021; Stuber et al. 2023), though remnants of a past bar and a possible nuclear ring have been debated (e.g., Sánchez-Blázquez et al. 2014a).
@@
M 74 is a star-forming galaxy, with a star-formation rate of $\log_{10}\,(\textrm{SFR}/[\textrm{M}_{\odot}/\textrm{yr}])=0.24$, which places it above the star-formation main sequence of local galaxies (Leroy et al. 2021; Kim et al. 2023). Its star formation and interstellar medium have been extensively studied at high spatial resolution by the PHANGS team using a combination of MUSE, ALMA, HST, and JWST. Massive star-forming regions are observed both in the spiral arm and interarm regions (Kreckel et al. 2016; Williams et al. 2022). On large scales, star formation mostly follows the molecular-gas structure, while offsets at the scale of molecular clouds imply longer gas depletion times than in the Milky Way (Kreckel et al. 2018). HST and JWST images revealed with unprecedented detail a complex web of dust and gas filaments, ranging from compact molecular-cloud-sized structures to faint arm-tracing features (Thilker et al. 2023), along with stellar-feedback-driven bubbles and voids (Barnes et al. 2023; Watkins et al. 2023). A remarkable cavity of a size of about $200\times 400$ pc2, with no ionized or molecular gas, and no dust, was found in the center of the galaxy (Kreckel et al. 2018; Herrera et al. 2020; Leroy et al. 2021; Stuber et al. 2023; Hoyer et al. 2023b).
@@
M 74’s spatially resolved stellar populations show radial trends. At large scale, Sánchez-Blázquez et al. (2014a) found predominantly old ages (${\sim}10$ Gyr) beyond ${\sim}6$ kpc, with a negative gradient consistent with an inside-out disk formation. The SFH reveals a dominant ${\sim}13$ Gyr component, plus a minor recent ${\sim}1$ Gyr component. The metallicities are slightly subsolar with a relatively flat gradient within ${\sim}3$ kpc, which becomes negative outwards. A more recent PHANGS-MUSE analysis (Pessa et al. 2023) confirmed these trends at ${\sim}100$ pc resolution, although with average light-weighted and mass-weighted ages in the central kpc, respectively, of about 2 and 9 Gyr, and the youngest stellar populations along the spiral arms. Metallicity gradients are negative in the central kpc and flatter (slightly positive) in the outskirts, again supporting an inside-out assembly scenario. JWST resolved-star analyses further confirmed the negative metallicity gradient (Ck et al. 2025). Finally, the star-cluster analysis from HST data has shown a distribution of ages ranging from a few million years, for cluster masses between $10^{3}$ and $10^{5}$ M, to more than 10 billion years above $10^{5}$ M(up to $10^{7}$ M, Whitmore et al. 2023).
@@
M 74 is the brightest member of the M 74 Group, in which dwarf satellite galaxies (with stellar masses $\rm 10^{4}\,M_{\sun}<M_{*}<10^{6}$ M) identified by Davis et al. (2021) show signs of recent quenching. M 74’s HI disk shows a complex structure with a warp, high velocity complexes, and an asymmetric outer tail (Kamphuis and Briggs 1992). These features might be the result of one or more past accretion events.